### v2

# The Potential of Citizen Platforms for Requirements Engineering of Large Socio-Technical Software Systems

Abstract
Participatory citizen platforms are innovative solutions to digitally better engage citizens in policy-making and deliberative democracy in general. Although these platforms have been used also in an engineering context, thus far, there is no existing work for connecting the platforms to requirements engineering. The present paper fills this notable gap. In addition to discussing the platforms in conjunction with requirements engineering, the paper elaborates potential advantages and disadvantages, thus paving the way for a future pilot study in a software engineering context. With these engineering tenets, the paper also contributes to the research of large socio-technical software systems in a public sector context, including their implementation and governance.
Keywords:
s
oftware engineering, large-scale software systems, civic engagement, deliberative democracy, public sector, participatory platforms
1 Introduction
In recent years there has been an increasing interest to better engage citizens in public policy-making due to real or perceived problems in representative democracy [1, 3, 14]. The argument is that citizens want to engage in politics and democracy, but they do not have knowledge or means to do so. The increasing digitalization of societies has allegedly also rendered the traditional participatory infrastructures outmoded; the old democratic agoras of a polity have disappeared or became dysfunctional. At the same time, large amounts of citizens have moved to social media platforms within which the many real or perceived problems are well-recognized; the list includes, but is not limited to, algorithmic polarization, engagement and amplification, so-called echo chambers and filter bubbles, disinformation, misinformation, and propaganda, hate speech and harassment, and many other issues [20]. According to critical viewpoints, deliberative democracy is in trouble due to these and other related trends [15].111 In addition to direct democracy, deliberative democracy is one of the leading concepts to patch the current problems in representative democracy. Drawing on ideas that can be traced all the way back to the ancient times, the essence is that “people come together, on the basis of equal status and mutual respect, to discuss the political issues they face and, on the basis of those discussions, decide on the policies that will then affect their lives” [1, p. 2]. To this end, these trends have prompted new arguments and ideas about the difficult relationship between democracy and its institutions, including knowledge, markets, and the so-called “marketplace of ideas”. For instance, it has been recently argued that expertise and epistemology are at the heart of the problems; citizens must have means to receive knowledge that matters for policy-making, and they must have functioning democratic channels to subsequently voice their knowledge-based opinions [9]. This argument suits well for framing also the paper’s scope.
Thus, on one hand, expertise and knowledge are essential also in requirements engineering; here, software or other engineers are the experts who solicit requirements from non-experts, including customers, clients, and stakeholders. To further frame the scope, the paper considers requirements engineering of large socio-technical systems in a public sector context; hence, the non-experts include also citizens for whom the systems are foremost implemented. Then, on the other hand, the previous points about platforms frame the paper’s central idea; the potential use of so-called citizen platforms for requirements engineering in the context described. These platforms have been successfully used in some countries for participatory policy-making and deliberative democracy in general.
While there is existing work also on the use of citizen platforms in large-scale engineering projects, such as so-called smart cities [13, 22], the paper is the first to contemplate these explicitly in relation to requirements engineering. The note about smart cities exemplifies what is meant by large socio-technical systems. Fundamentally, cities belong to the public sphere, and adding smart technologies to them thus necessitates the involvement of public sector stakeholders, including citizens, even in case the technologies themselves are designed, engineered, and operated by private sector companies. On these notes and clarifying framings, the opening Section 2 briefly elaborates what citizen platforms are about. The subsequent Section 3 continues by discussing their potential use in requirements engineering. A conclusion and a discussion are presented in the final Section 4.
2 Citizen Platforms
Citizen platforms are innovative means for better civic engagement in the current, increasingly digitalized societies. In essence: citizens gather on a platform for discussing, debating, and deliberating about a given public policy issue, eventually reaching a consensus everyone can agree upon. Taiwan has been the leading country in the adoption of citizen platforms; over the years, numerous platforms have been deployed there and successfully used to democratically solve policy issues [2]. Recently, citizen platforms have been developed or deployed also elsewhere, including in Finland [23] and Spain [3]. As a testimony of the potential offered, similar ideas and implementations have been adopted for even more challenging scenarios, such as peace-building and conflict resolution [4]. Against this brief background, it seems sensible that citizen platforms could be used also for requirements engineering in the public sector context in particular.
These platforms may overcome a long-standing sin in requirements engineering; the involvement of only a few stakeholders at the expense of users, citizens, employees, consumers, and other people for whom many systems are designed [6]. Given the public sector context, citizen platforms may also help at gathering voices from underrepresented or even marginalized citizen groups. The obvious benefit would be supposedly increased citizen satisfaction with a given socio-technical system implemented. In addition, deliberative participation may increase trust placed upon a system, among other things. Even more importantly, the platforms may catch two important birds with a single stone; not only may these increase satisfaction with and trust toward a socio-technical system, but these may also do the same with respect to democracy itself.
Three clarifications are needed before continuing. First: although the basic idea is rather similar, citizen platforms should not be confused with so-called citizen assemblies, which have been increasingly enacted in Europe in recent years [14]. Second: although somewhat similar in terms of technical implementations, citizen platforms are different from requirements engineering platforms designed for crowd-sourcing contexts [6, 8]. Importantly, citizens in a society are not a some anonymous crowd in the Internet; they have a right to participate and have their voices heard on public policy matters affecting them. The existing citizen platforms have also been strongly influenced by deliberation; participants on a platform seek a consensus through discussion. They are thus not merely a crowd from whom feedback is solicited. Third: from a broader perspective, citizen platforms also interlace with the discussion on public alternatives to the large multinational social media platforms and commercial platforms in general [25]. This point underlines the public sector context. It serves also as an argument in favor of citizen platforms; citizens and stakeholders might be more eager to participate on an open source citizen platform orchestrated by a public sector body due to privacy, data protection, integrity, trust, and ethical reasons.
In terms of existing implementations, Polis222 https://compdemocracy.org/Polis/ is the pioneering citizen platform. The discussions on the platform do not mimic those on social media or traditional discussion boards in the Internet. Instead, arguments are raised about a particular topic, and citizens can then agree or disagree with those. Participants can also raise new arguments for others to vote. Thanks to real-time analytics, participating citizens can also observe the formation of agreeing or disagreeing opinion groups. In contrast to the polarizing discussions and algorithms on social media, these features should help at reaching a consensus about a policy topic. Although the functionality may be slightly different in other implementations, such as Decidim [3]333 https://decidim.org/ and CONSUL DEMOCRACY444 https://consuldemocracy.org/, the basic ideas are still typically shared across different citizen platform implementations. Also tailored citizen platforms have been proposed [24]. Against this backdrop, it might be possible to develop a custom platform specifically for requirements engineering.
3 Citizen Platforms and Requirements Engineering
The functionalities of the Polis platform help at presenting a few high-level design ideas for such a platform. The iterative requirements engineering practices typically used in agile software development [18] can be used to frame the ideas. Thus, a requirements engineering process on a citizen platform might involve five phases: (1) a preparation phase, (2) an initial requirements engineering phase, (3) a prioritization phase, (4) a development phase, and (5) an acceptance phase.
Of these, the first phase would presumably be critical for successfully completing the later phases. Among other things, requirements engineers, possibly together with a product owner, would likely need to carefully write instructions and guidelines about a given large socio-technical system envisioned. These should probably also include high-level materials on matters such as user stories. Alternatively, it might be possible to arrange a preliminary offline or online lecture about these topics Some citizen platform implementations, such as the noted Decidim [3], have also been designed with education and pedagogy in mind. In addition to a potential risk of failing at delivering adequate instructions, the phase involves also a difficult question about how to recruit and invite citizens and stakeholders to a platform. It remains generally unclear how many citizens and stakeholders should be invited, and how representative they would be in terms of a whole society [21, 22]. While techniques from survey research might be used to gain representativeness for the recruitment invitations, it might well be that the resulting participants would be biased toward stakeholders on one hand and those citizens who are already engaged on the other. Such a democratic (digital) divide is well-known in the literature [17]. The potential bias would be particularly severe in case citizens for whom a system is geared are missing. For instance, it would seem oblivious to develop a public sector system for helping people to find employment in case unemployed people are missing.
In the second phase the citizens and stakeholders invited would engage on a platform by discussing a socio-technical system planned, raising high-level requirements as arguments, contemplating these with the help of analytics and a facilitating requirements engineer, and finally, if necessary, voting on the most important or preferable requirements. Also this phase as well as later gatherings on a platform contain their own risks.
As face-to-face requirements elicitation is preferred in agile development [18], digitalized platform-based elicitation would likely require new skills and techniques from software engineers. A similar point applies to participating citizens and stakeholders. Analogously to the so-called co-design practices, it might be that participants with technological or domain-specific knowledge would dominate those without such knowledge; so-called end-users would be overruled by so-called power users who are closer to being developers themselves [7, 21]. A similar point applies with respect to powerful stakeholders and interest groups.
It might also be that ordinary citizens would have difficulties to articulate requirements. This point underlines the importance of the prior instructions and guidelines. Analogously to focus groups [12], the role of a facilitating and moderating requirements engineer is likely highly important too [3, 6]. If there are difficulties to generate requirements, or if a discussion gets stuck, he or she should have the skills to put the process again on the right track. Given the public sector context, it might also be that a discussion gets too heated; hence, he or she should also know how to moderate a discussion, possibly by using a platform’s moderation tools. Because citizen platforms should engage also citizens who possibly speak different languages [22], language barriers might pose an additional risk, including with respect to a facilitator.
After the initial gathering on a platform, requirements engineers and other developers would engage during the prioritization phase by evaluating particularly the feasibility of the requirements put forward, ranking these, and sketching design ideas for implementing these. Given the context of large public sector socio-technical systems, the evaluation process should involve also public sector representatives with expertise in financial, legal, and related matters. Then, the prioritized requirements would be either delivered to the citizens and stakeholders directly or a second gathering would be arranged on a platform for the citizens and stakeholders to evaluate the prioritized requirements, possibly again voting on these and their ranking. There are risks affecting also this phase, at least in the form sketched. Because civil servants would be likely needed in the feasibility evaluations, the process sketched would imply a double burden for requirements engineering. In addition, it might be that not the citizens and stakeholders on a platform but the civil servants would fail to reach a consensus. Particularly financial matters might be difficult to agree upon. A further point is that legal requirements are not easy to specify [10]. Given the public sector context, financial uncertainties and legal obstacles might even lead to a failure of the prioritization phase altogether. To counter these risks, it might be possible to use a different platform instance for this auxiliary group of participants; after all, they are—or should be—experts in deliberation already. Some citizen platform implementations [3] offer also tools for budgeting and project financing.
The development and acceptance phases would then proceed as is common in agile software development. After an iteration of development, a gathering might again be arranged on a platform for the citizens and stakeholders to evaluate and discuss a given system being developed. Either voting or informal consensus-seeking might be again used to evaluate satisfaction with the system together with possibly new requirements that may emerge in the discussions. Given that the amount of participating citizens and stakeholders would presumably be much larger than in conventional software projects, it might be reasonable that a single gathering on a platform would cover multiple development iterations. Another, alternative option might be to impose fixed time limits for the deliberations. Regardless, during the final acceptance phase a finished system would be evaluated particularly against the initial and subsequent requirements. While voting or informal deliberation might be again used, it might be also possible to develop a citizen platform in a way that supports the ideas of acceptance testing.
Furthermore, it may be that plain voting, while necessary for a democracy, is not particularly well-suited for requirements engineering, despite it sometimes appearing in the literature [8, 19]. This point applies to co-design and co-creation as well. A memorable illustration of the point appeared in the Simpsons cartoon series; in an episode Homer had a chance to design a new car, and, expectedly, it failed miserably because a lot of unneeded and undesirable features were implemented.555 https://simpsons.fandom.com/wiki/The_Homer In other words, so-called feature creep too is a risk for large-scale participatory requirements engineering on a citizen platform. Given that deliberation and consensus-seeking are continuous endeavors, the risk further entails the commonplace problem of changing requirements, which have often been seen as a major cause for feature creep [11]. Due to these and related reasons, it may also be that citizen platforms are a poor solution to develop innovative systems. Though, unlike in product development, it must be added that innovativeness is not necessarily a primary or even a top-ranked criterion for the development of large socio-technical systems in the public sector context.
Against these backdrops, nevertheless, a citizen platform for requirements engineering may need more complex voting techniques that take different feasibility and related criteria into account, including financial, legal, and other constraints that are important for requirements engineering and software design [5, 10]. It might also be reasonable to prototype and evaluate so-called downvoting functionality. Even with such techniques and functionalities, a key point should be kept in mind: the “better the epistemic processes of deliberation before voting—including, where needed, the transmission of relevant expert knowledge to the broader public—the more likely it is that votes will be epistemically efficient in the sense that individuals are well informed about the subject matters at stake” [9, p. 71]. A further important point about voting is related to security and privacy. In case sensitive matters are voted upon, confidentiality should be guaranteed analogously to electoral secrecy in conventional democratic elections.
As has been also pointed out in a related context [8], a citizen platform for requirements engineering may also benefit from customized algorithms, including natural language processing methods tailored for requirements specifically. These methods are also a thriving research domain in requirements engineering [26]. Additionally, techniques such as gamification might be used [6], although the so-called nudging involved in these techniques arguably aligns poorly with the fundamental ideals of deliberative democracy. In other words, even in the presence of tools and algorithms, the central design idea should be that citizens and stakeholders can freely, without manipulation or other push factors, reach a consensus about requirements in their own terms through informed discussions. For such deliberation to work properly, openness and transparency are particularly important factors in the public sector context [17]. A lack of these has often been seen as a major problem for the today’s large commercial social media platforms [20]. To this end, it suffices to end the discussion by emphasizing a risk of replicating existing problems insofar as deliberative democracy is considered.
4 Conclusion
This short vision paper presented preliminary but novel ideas about a potential use of citizen platforms for requirements engineering of large public sector socio-technical systems. In contrast to the crowd-sourcing ideas already presented in the requirements engineering literature, these platforms are essentially about consensus-seeking through deliberation, including on potentially controversial or otherwise hot political topics.
By and large, these platforms have been previously used for politics and policies, not implementations based on the policies enacted. In terms of governance of large public sector socio-technical systems, both citizens and stakeholders exert pressure upon political decision-makers who enact new technology regulations [16]. By implication, in order to reach the full potential of civic engagement, citizen platforms could (or should) be used for both the formulation of technology policies and the implementation of conforming software solutions. There is a lot to improve also on the traditional policy side. For instance, in the European Union (EU), from which most technology policies nowadays come from in Europe, the associated policy consultations can gather thousands of responses [20], and it remains unclear how well the voices of EU citizens are heard in the face of heavy lobbying by powerful industry groups. In addition, policy consultations have obvious limitations in that these are largely non-deliberative and particularly non-iterative. Thus, maybe there might be room to pilot citizen platforms also in the EU’s future technology policy-making. Having said, it remains generally unclear how the potential risk of hijacking or otherwise dominating a deliberation process could be addressed. This risk reinforces the earlier points raised about representativeness, facilitation, knowledge-delivery and epistemology in general, and even content moderation.
Piloting a citizen platform for requirements engineering would be a good way to continue in research. A pilot study should shed more light on the potential advantages and disadvantages, including the risks discussed. In particular, better understanding is required on whether and how well the principles of deliberative democracy apply to requirements engineering. In this regard, there are two fundamental evaluation dimensions to consider. The first is the outcome-side [17], that is, in the present context the quality and other related properties of a socio-technical system designed and implemented. Such an evaluation requires that a pilot should involve a design of a real-world system. The second is the platform-side, including with respect to the risk discussed. On this side, an evaluation should try to also reflect upon a risk of potential participatory washing, meaning the hiding of existing problems in democracies underneath platforms and technology in general [3]. Then, if a pilot study indicates that citizen platforms are poorly suited for requirements engineering, it might be reasonable to contemplate alternatives. An example would be public procurement of large socio-technical systems; here, a citizen platform might be used to evaluate tendering in terms of acceptability and desirability among citizens. As was discussed, a further research avenue opens on the implementation side; it might be desirable to implement a new citizen platform specifically for requirements engineering. Though, in order to avoid wasting resources, the applicability of the existing citizen platform implementations should be evaluated first.
Disclosure of Interests
There are neither competing nor conflicting interests.